convert 46 to base two?
convert 110101

Answer by solver91311(24713)


46 < 64
46 - 32 = 14
14 < 16
14 - 8 = 6
6 - 4 = 2
2 - 2 = 0
0 < 1

One 32
Zero 16s
One 8
One 4
One 2
Zero 1s

101110

+---------+---------+---------+---------+---------+---------+

110101

32 + 16 + 4 + 1 = 53



John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism